1 In a competitive market, the industry demand and supply curves are P= 200 - .2Qd and P = 100 + .3Qs, respectively. Find the market’s equilibrium price and output
1
Expert's answer
2017-02-28T11:18:05-0500
If in a competitive market, the industry demand and supply curves are P = 200 - 0.2Qd and P = 100 + 0.3Qs, respectively, then the market’s equilibrium price and output are in the point, where Qd = QS, so:
200 - 0.2Q = 100 + 0.3Q,
0.5Q = 100,
Q = 200 units.
P = 200 - 0.2*100 = $180.